OpenStudy (anonymous):

Consider the reaction of aluminum with hydrogen chloride to produce aluminum chloride and hydrogen gas. If you react 151g of aluminum with 592g of hydrogen chloride, how much hydrogen gas could be produced, in grams?

OpenStudy (anonymous):

balanced equation.

OpenStudy (anonymous):

The balanced equation is 2 Al + 6 HCl = 2 AlCl3 + 3 H2

OpenStudy (anonymous):

The equal sign shud be a yield sign

OpenStudy (anonymous):

I found the moles of A and moles of HCl but im not sure if that's necessary

OpenStudy (anonymous):

i think the answer is 16.9 because the HCl is the limiting reactant right?

OpenStudy (anonymous):

16.9g H2

OpenStudy (anonymous):

write the balanced eqn. find the Limiting reactant.

OpenStudy (anonymous):

convert hcl to moles then mole mole ratio then mass h2

OpenStudy (anonymous):

Aluminum is the limiting reagent i think since its 0.55 mol and HCl is 16.24 mol

OpenStudy (anonymous):

How did u get 16.9. like what were ur steps

OpenStudy (anonymous):

16.4 sorry

OpenStudy (anonymous):

yeah that is teh right answer b/c i have teh asnwer key but i dont understand how to get taht answer

OpenStudy (anonymous):

well i determined the theoretical yield of h2 produced from the balanced eqn

OpenStudy (anonymous):

and you to that by

OpenStudy (anonymous):

taking 151g Al, convert to moles, (times)mole ratio of aluminum to h2, then multiply by the mass of h2

OpenStudy (anonymous):

and that will give you the mass of h2 produced.

OpenStudy (anonymous):

but you also have 592g HCl so u do the same exact thing except with HCl

OpenStudy (anonymous):

THE SMALLER NUMBER IS THE THEORETICAL YIELD or the amount that can be produced by the reaction

OpenStudy (anonymous):

theoretically

OpenStudy (anonymous):

so i got 0.55 mol Al. Now I multiple it by (3 H2/ 2 Al) = 0.825 mol H2 and 16.24 mol HCl 8 (3 mol h2/ 6 mo HCl) = 8.12 mol HCl.

OpenStudy (anonymous):

ok reset bro start with grams Al

OpenStudy (anonymous):

i mean HCl is limiting reagent right? then u multiply 8.12 * 2 = 16.24 g

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

8.2 is the number of moles of hydrogen

OpenStudy (anonymous):

ok just alittle confusion. Why is the larger number of moles the limiting reagent?

OpenStudy (anonymous):

because mole mole ratio changes it

OpenStudy (anonymous):

Oh I get it now! Thanks :)

OpenStudy (anonymous):

yeah so for this problem start with grams, goto moles, and back to grams. also make sure you keep track of units

OpenStudy (anonymous):

units are very important

OpenStudy (anonymous):

youre welcome btw. take a look at the question i posted if u know where to start

OpenStudy (anonymous):

lmk

OpenStudy (anonymous):

ok sure